OpenStudy (anonymous):

2 log x = -1 solve the equation ??

OpenStudy (anonymous):

First let's isolate log(x) 2 log(x) = -1 log(x) = -1/2 When you have log just by itself with no specific base, we assume it's base 10. So bascially what is the question asking you? Well isn't just asking you 10^(-1/2) = x x = 10^(-1/2) x = 1/[10^(1/2)] \[x = 1/(\sqrt{10})\]